Suppose Y=Wx²/6Ei(3a-x) for x<=a
Y=Wx²/6Ei(3x-a) for x>a
Use the if function to compute the value of deflection if W is in cell A4,E in A5,I in A6,a in B6 And x in B8
1
Expert's answer
2020-02-17T15:35:41-0500

Final formula

=IF(B8>B6;((A4*B8^2)/(6*A5*A6))*(3*B8-B6);((A4*B8^2)/(6*A5*A6))*(3*B6-B8))


Explanation

We have variables located this way:




IF formula consists of 3 parts: condition, expression if condition is true, expression if condition is false.


Condition: we would check if x>a

Excel representation: B8>B6


If condition is true:

Formula for x>a is "Y=\frac{Wx^2}{6Ei}(3x-a)"

Excel representation: ((A4*B8^2)/(6*A5*A6))*(3*B8-B6)


If condition is false:

Formula for x<=a is "Y=\frac{Wx^2}{6Ei}(3a-x
)"

Excel representation: ((A4*B8^2)/(6*A5*A6))*(3*B6-B8)


Combining theese 3 parts we would receive final formula

=IF(B8>B6;((A4*B8^2)/(6*A5*A6))*(3*B8-B6);((A4*B8^2)/(6*A5*A6))*(3*B6-B8))